2+1=5 3+2=11 4+3=19 5+4=29 6+7=?? only for genius.................................................................................

2+1=5
3+2=11
4+3=19
5+4=29
6+7=??
only for genius.................................................................................

Grade:11

3 Answers

SAHIL
6 years ago
HI 
ANSWER IS 53
BECAUSE 2 X 1 + 3 = 5
ALSO 3 X 2 + 5 = 7
ALSO 4 X 3 + 7 = 19
SAME AS 5 X 4 + 9 = 29
SO ANSWER IS 6 X 7 + 11 = 53
THANKS AGAIN
NAVEEN PRAVEEN
6 years ago
HI MY DEAR FRIEND ….….
ANSWER IS 2×2+1=5
                   3*3+2=11
                   4*4+3=19
                   5*5+4=29
                   6*6+7=43....................
      THE ANSWER IS 43  
 
imran
6 years ago
Hi  CHOWDAIAH
I think answer is 43
because by observing the above conditions 
i can say that the first digit is multiplied by itself to get the answer
                                              2*2+1=5
                                              3*3+2=11
                                              4*4+3=16
                                              5*5+4=29
                                   in that way we can say that 6*6+7=43
       OK THANK Q